OpenStudy (anonymous):

How to solve for x? 11.04=10(1.02)^x

OpenStudy (phi):

I would first divide both sides by 10.

OpenStudy (anonymous):

So, then I have 1.104= (1.02)^x right?

OpenStudy (phi):

yes. now "take the log" of both sides: log(stuff on left) = log(stuff on right)

OpenStudy (phi):

we can assume log base 10 on the right side you can use this "rule": \[ \log(a^b)= b\log(a) \]

OpenStudy (phi):

\[ \log\left(1.104 \right)= \log\left(1.02^x \right) \]

OpenStudy (anonymous):

That's kind of over my head... I'm not sure what the "log" is?

OpenStudy (phi):

You need to use logarithms to solve it. You must have studied this. If you need a refresher, see https://www.khanacademy.org/math/algebra2/logarithms-tutorial/logarithm_basics/v/logarithms

OpenStudy (phi):

and here is an example of solving this type of equation https://www.khanacademy.org/math/algebra2/logarithms-tutorial/logarithm_basics/v/exponential-equation though in your problem, you will need a calculator.

OpenStudy (anonymous):

Okay so at this point, I have to solve log1.02 1.104=x ?

OpenStudy (anonymous):

Which is 1.02^5?

OpenStudy (phi):

how did you get that ?

OpenStudy (anonymous):

It's what they did in that first khan academy video, and I just substituted my numbers in...

OpenStudy (phi):

\[ \log\left(1.104 \right)= \log\left(1.02^x \right) \] re-write the right side using \( \log(a^b) = b \log(a) \) can you do that ?

OpenStudy (phi):

here is an example with just numbers: log(10) = 1 (try it with a calculator) log(10^2) = 2 log(10) = 2*1 = 2 as a test, we know 10^2 = 10*10 = 100 and log(100)= 2 the rule works!

OpenStudy (phi):

in your problem match the "a" with 1.02 and "b" with x

OpenStudy (anonymous):

I'm sorry I'm not quite getting what you're telling me, but that khan academy link you sent really helped. I got the answer right... Thanks for helping!